OpenStudy (anonymous):

1. What is the quotient x^2-16/2x^2-9x+4 divide by 2x^2+14x+24/4x+4 2What is the simplified form for the product x+1/x^2-25 time x+5/x^2+8x+7 3.What are the coordinates of the x-intercepts of the graph of y = 2x^2 + 6x – 20? 4.What is the simplified form of the product x^2+8x+15/x-4 times x^2-16/2x+6

OpenStudy (campbell_st):

factorise the quadratics where possible look at the numerator 1st \[\frac{(x-4)(x+4)}{(2x-1)(x -4)} = \frac{(x +4)}{(2x-1)}\] now for the denominator \[\frac{2(x^2 + 7x + 12)}{4(x +1)} = \frac{(x +3)(x+4)}{2(x +1)}\] dividing by a fraction requires the denominator to be flipped and then multiplied \[\frac{(x +4)}{(2x -1)} \times \frac{2(x+1)}{(x+3)(x+4)}\] cancel the common factor (x +4) leaves \[\frac{2(x +1)}{(2x-1)(x+3)}\] all you need to do is expand and collect like terms \[\frac{2x + 2}{2x^2 + 5x - 3}\]

OpenStudy (anonymous):

1) C 2) C 3) A 4) B these are correct!